OpenStudy (jbrocato22):

A ball is thrown vertically upward from the top of a 100 foot tower, with an initial velocity of 10 ft/sec. Its position function is s(t) = –16t2 + 10t + 100. What is its velocity in ft/sec when t = 2 seconds?

OpenStudy (irishboy123):

are you learning calculus? if so, compute \(\frac{ds}{dt}\) at t = 2 if not, you are using Imperial measurements (yuk), and so the thing has velocity +10 and as you can see from the equation acceleration -32 so the equations of motion should get you there too.....

OpenStudy (jbrocato22):

yeah I'm in calculus and I'm having a tough time figuring this one out

OpenStudy (shamim):

U know Velocity v=ds/dt=d/dt(-16t^2+10t+100)=-d/dt(16t^2)+d/dt(10t)+d/dt(100)=?

OpenStudy (shamim):

=-16*2t+10+0=-32t+10

OpenStudy (shamim):

So velocity v=-32t+10

OpenStudy (shamim):

Now just put t=2 sec in my last equation

OpenStudy (jbrocato22):

thank you so much!